4.

Janus Corporation was unable to find a store suitable for its business, so it decided to build one. It was able to secure debt financing from the Southeast Bank in the amount of $4,000,000 at an interest rate of 5 percent. During 20X8, Janus spent $2,500,000 on construction, but did not complete the building. Janus continued work on the building into 20X9, eventually completing it on July 1 at a total cost of $3,800,000. Janus does not use the half-year convention.

1.

Determine the amount at which Janus should record the building, including any applicable capitalized interest.

2.

If Janus expects the building to be in use for twenty years with negligible residual value, what would depreciation expense be in 20X9?


1
Expert's answer
2021-07-20T10:47:09-0400

1.

The amount to be capitalized is equal to Interest on loan till the building is ready to use + Total Cost incurred


Here total interest till construction = interest for 2018 + half year interest of 2019


Interest = $4000000 * 5% + $4000000 * 5% * "\frac{6}{12}" = $300000


The Total amount to be capitalized = $3800000 + $300000 = $4100000


2)

The Depreciation will be Charged for the whole year of 2019 as Janus Corp. does not use a half-year convention


 So Depreciation = "( cost - residual value ) \/ life of asset"

 here Depreciation =


               "\frac{(4100000 - 0)}{20}" years = $205000